How many molecules of sodium oxide will be created if 275 grams of sodium reacts with excess oxygen?

Given the equation: 4Na + O2 -> 2Na2O

A.

72 x 10231023 molecules of Na2ONa2O

  • B.

    108 x 10231023 molecules of Na2ONa2O

  • C.

    12.04 x 10231023 molecules of Na2ONa2O

  • D.

    36 x 10231023 molecules of
Answer #1

Given

275 grams of sodium reacts with excess of oxygen

no of moles of sodium = mass / molar mass

we know that,

molar mass of sodium = 23 g/mol

no of moles of sodium = 275 g / 23 g/mol = 11.96 moles

Given the equation:

4Na + O2 -> 2Na2O

from the equation,

If 4 mole of sodium react, then 2 mole of Na2O will result.

for 11.96 moles ==> 11.96 / 2 = 5.98 moles of Na2O will result

we know that

1 mole = 6.022×1023 molecules (Avogadro's number).

hence

5.98 moles of Na2O = 5.98 * 6.022×1023 molecules of Na2O = 3.6 * 1024 molecules of Na2O

   = 36 * 1023 molecules of Na2O